OpenStudy (anonymous):

The average age of vehicles registered in the United States is 8 years, or 96 months. If a random sample of 36 vehicles is selected, find the probability that the mean of their age is between 98 and 100 months? Assume the standard deviation for the population is 15. Choose one answer. a. 63.47% b. 75.82% c. 66.67% d. 15.71%

OpenStudy (amistre64):

it looks like we find 2 zscores and find the area between them: (98-96)/15 = z1 = .1333 (100-96)/15 = z2=.2666

OpenStudy (amistre64):

z1 gives me .0517 z2 gives me .1026 z2-z1 should be the area between them: .1026 .0517 ------ .0509 .... i forgot to modify the sd :) 15/sqrt(36) is what we needed

OpenStudy (amistre64):

(98-96)(sqrt(36)/15 = z1 = .8 (100-96)(sqrt(36)/15 = z2= 1.6 z1 = .2881 z2 = .4452 .4452 .2881 ------ .1571 = 15.71% :)